# Quasistatic approximation in neuromodulation

### v4

@@ Abstract
We define and explain the quasistatic approximation (QSA) as applied to field modeling for electrical and magnetic stimulation. Neuromodulation analysis pipelines include discrete stages, and QSA is applied specifically when calculating the electric and magnetic fields generated in tissues by a given stimulation dose. QSA simplifies the modeling equations to support tractable analysis, enhanced understanding, and computational efficiency. The application of QSA in neuromodulation is based on four underlying assumptions: (A1) no wave propagation or self-induction in tissue, (A2) linear tissue properties, (A3) purely resistive tissue, and (A4) non-dispersive tissue. As a consequence of these assumptions, each tissue is assigned a fixed conductivity, and the simplified equations (e.g., Laplace’s equation) are solved for the spatial distribution of the field, which is separated from the field’s temporal waveform. Recognizing that electrical tissue properties may be more complex, we explain how QSA can be embedded in parallel or iterative pipelines to model frequency dependence or nonlinearity of conductivity. We survey the history and validity of QSA across specific applications, such as microstimulation, deep brain stimulation, spinal cord stimulation, transcranial electrical stimulation, and transcranial magnetic stimulation. The precise definition and explanation of QSA in neuromodulation are essential for rigor when using QSA models or testing their limits.
@@ 1 Introduction
Computational modeling of electromagnetic fields has broad application in neuromodulation to estimate delivered intensity, predict and optimize response, and inform risk. The quasistatic approximation (QSA) [1, 2, 3, 4, 5, 6, 7], applied under the associated quasistatic assumptions, is ubiquitous in computational modeling of electrical and magnetic stimulation. The purpose of QSA is to reduce the computational complexity of calculating the electric fields (E-fields), potentials, and currents in the target tissue, e.g., brain, spinal cord, or peripheral nerve. Despite its prevalence, QSA as used in neuromodulation has not been clearly defined, and the underlying assumptions are typically implicit or, at times, ambiguous. Thus, the use of QSA, as understood and applied in neuromodulation modeling, needs to be defined. Here we explain how QSA is applied in a consistent manner across modeling of neuromodulation under four underlying assumptions, and we derive the corresponding governing equations. We also review adaptations of the QSA simulation pipeline to accommodate cases where some of the simplifying assumptions do not hold. We further provide historical and application-specific examples and clarify QSA-related terminology as applied in general electromagnetism. Finally, we provide recommendations for communicating the use of QSA in modeling studies.
@@ 1.1 Scope
The scope of this paper includes the computational methods used to predict the delivery of electromagnetic energy to the body during neuromodulation—and specifically the application of QSA. These computations yield the E-fields in the tissue of interest, and/or related quantities including tissue potentials, current densities, or magnetic fields. The solutions provide estimates of these variables, which are subject to model assumptions and verification. We define these assumptions and related parameters, focusing on the E-field, as the magnetic field is not influenced by the tissues when QSA is valid, and as the E-field is understood to mediate the effects of the most common forms of neural stimulation [8]. Nonetheless, the magnetic field has a key role in neural stimulation through electromagnetic induction, and QSA typically applies to the magnetic field as well; therefore, we occasionally note generalization to the magnetic field. To frame the use of QSA, we also broadly consider additional steps in modeling pipelines, before or following the prediction of the E-field. Quasistatic analysis of approaches for recording bioelectric phenomena, such as the electrocardiogram [9] and impedance plethysmography, was summarized by Plonsey and Heppner [1]. Specific to the nervous system, QSA has been applied to analyses of endogenous brain activity as recorded by electroencephalography (EEG) [10, 11] and magnetoencephalography (MEG) [10, 12], as well as recording of local field potentials or evoked potentials using (micro)electrodes in the cortex or from peripheral nerves [13, 14]. Subsequently, QSA in neuromodulation was adapted from these analyses of bioelectric signal recording. In contrast to endogenous electromagnetic fields, however, neuromodulation uses exogenous fields and applies a wide range of stimulation waveforms with much broader frequency content and much higher intensities than intrinsic signals, and thus warrants distinct considerations. The application of QSA in modeling neuromodulation encompasses several modalities, including:
@@ 2.1 Multi-stage framework for modeling of neuromodulation
In some modeling pipelines, components of the stimulation device (electrodes or coils) are combined with the tissue when the E-field is computed. This combination preserves QSA as long as all underlying assumptions applied to the tissue also apply to the device components. Alternatively, the influence of the electrodes and their interface with tissue can be modeled in a step after defining the current or voltage source and preceding tissue E-field calculation [24, 25], in which case the electrodes are not subject to QSA. For example, a lumped-circuit model can represent the impedance of the electrode–tissue interface. Similarly, the magnetic field generated by a TMS coil can be computed prior to simulation of the induced E-field in tissue, and can account, for example, for nonlinear effects due to saturation of a ferromagnetic core in the coil [26]. These approaches preserve QSA for the E-field in the tissue, allowing for modeling adaptations (section 2.6) to deal with sources with frequency-dependent or nonlinear effects, such as charge transfer conductances of the electrode–electrolyte interface in electrical stimulation and skin effect in coil windings or saturation of ferromagnetic core materials in magnetic stimulation.
@@ 2.2.3 Purely resistive tissue (A3)
Under assumption A3, the time derivative relating the current density and E-field is neglected. Only the free current density and thus the real part of the complex conductivity is considered. This removes the frequency- and location-dependent temporal phase shift between the E-field and current density distributions due to the tissue capacitance. This assumption is typically valid for low permittivity media at low frequencies, and is commonly denoted as “no capacitance” or a “purely resistive” medium in the neuromodulation modeling literature.
@@ 2.2.4 Non-dispersive tissue (A4)
Assumption A4 considers tissue properties to be non-dispersive, indicating that they remain frequency-independent, at least within the spectrum of the stimulation waveform. In practice, neuromodulation modeling studies typically make this assumption implicitly, along with assumptions A2 and A3, by assigning a single conductivity value to each tissue. For stimulation using a single frequency (e.g., tDCS or tACS), these tissue parameters reflect the magnitude of the impedance at the given frequency. In these cases, assumption A4 is not so much an assumption of frequency-independent tissue properties as the (implicit) recognition that the relevant properties correspond to a specific frequency. For the majority of cases (e.g., pulsed stimulation), waveforms are composed of multiple frequencies (see section 2.4 and Fig. 2 A), and the use of a single parameter for each tissue reflects an approximation. Such an approximation may be justified by considering the central frequency content of the waveform, or otherwise sufficient given the goals of the stimulation study. In all cases, the selection of single fixed purely-resistive tissue values (which involves assumptions A1, A2, and A3) should be recognized as a key step in model development (see section 4).
@@ 2.3.1 Current density in tissue volume conductor
We start by deriving the equation that governs the electric current density $\boldsymbol{J}$ in the tissue under the generalized physical case. The current density, absent magnetizing currents in a non-magnetic volume conductor, consists of the free current and displacement current (including the polarization current) driven by the position-dependent and time-varying E-field, $\boldsymbol{E}({\color[rgb]{0,0,0}\boldsymbol{r},}t)$,
@@
Current continuity on the boundary of two tissue domains 1 and 2 [64, 65] results in $\sigma_{1}\cdot\boldsymbol{E_{1,\perp}}=\sigma_{2}\cdot\boldsymbol{E_{2,\perp}}$ for the normal component, or specifically $\boldsymbol{E_{1,\perp}}=0$ if domain 2 is non-conductive, such as air or electrode insulation. In general, the primary E-field impressed by the stimulation sources (i.e., electrode charges or coil currents) does not satisfy (3) and the boundary conditions. The resultant imbalance of current density therefore induces charge densities on tissue boundaries, in a process considered instantaneous under QSA [9, 63]. For anisotropic or isotropic but not piecewise homogeneous medium, a volumetric charge density will additionally appear and contribute to the secondary field [66]. The conservative secondary field by the induced tissue charges ensures that the total E-field complies with (3) and the boundary conditions under QSA.
@@ 2.3.2 Electrical stimulation
For electrical stimulation, both the primary and secondary E-fields are conservative and therefore typically not distinguished. Defining the total E-field in (3) as the gradient of a scalar potential $\varphi$, $\boldsymbol{E}=-\nabla\varphi$, results in the governing equation for electric potential with the real-valued, fixed, and frequency-independent conductivity, which is commonly used in neuromodulation:
@@ 2.3.3 Magnetic stimulation
where $\boldsymbol{A}$ is the magnetic vector potential that is related to the coil magnetic field $\boldsymbol{B}$ by $\boldsymbol{B}=\nabla\times\boldsymbol{A}$. Here, despite the temporal variation being considered very slow (relative to typical radio frequencies [69]) under QSA, the primary field $\boldsymbol{E_{\mathrm{p}}}=-\frac{\partial\boldsymbol{A}}{\partial t}$ generated by the forward mutual induction between the coil and tissue is included because the rate-of-change is high due to the very high coil currents (on the order of kiloamperes). In contrast, the induced currents in the tissue (less than 1 A) are several orders of magnitude smaller than the kA-level coil currents. Therefore, their magnetic field and induced E-field—both self-induction in the tissue and backward mutual induction to the coil—are negligible. The secondary field $\boldsymbol{E_{\mathrm{s}}}=-\nabla\varphi$ from the induced tissue charges needs to be calculated separately. Given (3), the scalar potential is related to the vector potential by [70, 71, 72, 4, 73]
$\nabla\cdot(\sigma\nabla\varphi)=-\nabla\cdot(\sigma\frac{\partial\boldsymbol{A}}{\partial t})=\nabla\cdot(\sigma\boldsymbol{E_{\mathrm{p}}})$ (7)
and can be obtained numerically via techniques such as the finite- or boundary-element method (FEM or BEM, respectively), with the latter specifically representing and utilizing the boundary charges [9, 71, 4, 73, 74, 75, 76, 77].
In magnetic stimulation, the presence of the volume conductor comprising tissues does not significantly affect the electrical characteristics of the coils or their magnetic field. Therefore, the coil field can be pre-calculated or measured in the form of the magnetic field ($\boldsymbol{B}$ or $\boldsymbol{A}$) or the primary field $\boldsymbol{E}_{\mathrm{p}}$ and applied as a background field in the numeric solver [72, 26]. The field solution is computed for either a specific rate of change of the coil current (e.g., 1 A/µs) or a value to match the coil current waveform characteristics. Hence, the electromagnetic induction is handled in a stage preceding the quasistatic modeling of the E-field. Alternatively, the coil(s) can be explicitly included as part of the tissue E-field model, and a frequency-domain simulation may be used to calculate the coil’s magnetic field and induced E-field with the appropriate current amplitude and frequency to reflect the coil current’s rate of change [78].
@@ 2.4.1 Frequency spectrum of common stimulation pulses
tDCS applies constant currents via scalp electrodes (disregarding slow ramp-ups and ramp-downs). Usually, tACS is applied using only a single frequency. When the stimulation waveform is a relatively high frequency sinusoid that is amplitude modulated, the frequency spectrum is centered around the carrier frequency—not around the lower frequency of the amplitude modulation. This is also the case for TIS, where the frequency spectrum consists of the relatively higher frequency of the two (or more) individual channels and does not contain low frequency energy [79, 80, 81].
@@
The frequency spectrum of pulsed stimulation depends on the pulse width, shape, and repetition patterns. While monophasic pulses cover a wide frequency range starting from 0 Hz (Fig. 2 Ai), neural stimulation almost always uses charge balanced stimulation pulses (Fig. 2 Aii–Aiii) that have energy typically distributed within a frequency range of around one decade. The center location of the spectrum is determined by the pulse duration, typically falling between 100 Hz and 100 kHz. Increasing the total pulse duration–whether by increasing the phase duration, including an interphase interval, or using an asymmetric pulse with a longer and lower amplitude secondary phase–reduces the center frequency of the spectrum but has smaller effects on its shape. The pattern of pulse stimulation (e.g., continuous versus burst) also does not typically change the overall range or shape of the spectrum as governed by the individual pulses, but rather results in a discretized spectrum with the sampling and envelope determined by the repetition pattern and total duration of the pulse train. The spectra of the stimulation waveform should be considered when selecting the conductivity of tissue(s) within neuromodulation models.
@@
Figure 2: A. Normalized power spectrum density between 0 and 100 kHz of representative neurostimulation pulse waveforms, calculated using a 1 µs sampling rate and a total time interval of 1 s for single pulses. i. Monophasic pulses with pulse width (PW) of 1 ms (blue: single pulse; light blue: 30 s pulse train at 20 Hz) and 100 µs (red: single pulse). ii. Charge-balanced biphasic pulses with a first phase of 100 µs duration and with a second phase of equal duration (blue: single pulse; light blue: 15 s pulse train at 125 Hz), a second phase of equal duration and with 100 µs interphase/intrapulse interval (IPI) (red: single pulse), and second phase with 500 µs duration and 20% amplitude of first phase (green: single pulse). iii. E-field waveforms induced by magnetic stimulation, including monophasic (blue: single pulse) and biphasic (red: single pulse; orange: continuous theta burst stimulation (cTBS), 3 pulses per burst, 50 Hz intraburst frequency, 5 Hz interburst frequency, and 40 s duration) pulses, recorded from a MagPro X100 device with a Mag-Venture MCF-B70 figure-of-8 coil (MagVenture A/S, Farum, Denmark) [36, 41]. Electromagnetically-induced pulses are automatically charged-balanced as the amplitude of the coil current returns to zero at the end of a pulse. B. Spectra of tissue electrical parameters between 0 and 1 MHz for gray matter (blue) and white matter (red) as modeled by Gabriel et al. [60]. i. Relative permittivity. ii. Resistive (darker colors) and capacitive (lighter colors) conductivities. iii. Ratio between conductive and capacitive conductivities. Conductivity variations are small over the frequency ranges of charge-balanced pulsed stimulation, enabling the approximation using single conductivity values. Dashed lines indicate extrapolation of the model to very low frequencies ($<10$ Hz), outside the range of experimental data (10 Hz to 20 GHz) to which the model was fitted and where the model has unknown validity. Due to the experimental setup, the measurements also had high uncertainties in the low frequency range (10 to 100 Hz), for which updates from other sources have been suggested by the IT’IS Foundation database [82]. All horizontal axes use a mixed scale, with linear scale between 0 and 10 Hz and logarithmic scale above 10 Hz.
2.4.2 Frequency dependence of tissue electric parameters and influence on QSA accuracy
@@
In 1996, Gabriel et al. published a series of papers on the electrical conductivity and permittivity of tissue at frequencies from 10 Hz up to 20 GHz [83, 84, 60]. This series was subsequently updated [85] and combined with other low-frequency data sources in the IT’IS Foundation database [82]. Their data and models, which fit particularly well at high frequencies, are broadly consistent with the dielectric properties measured in other studies and as described by the Cole–Cole equation, and have, to some extent, become a canonical reference for modeling studies explicitly considering the suitability of QSA.
@@
According to the data by Gabriel et al., the capacitive current of neural tissues in the relevant frequency range is $\sim 10\%$ compared to the resistive current (Fig.2 B). The total current density consists of the resistive and capacitive currents in parallel and is driven by the same E-field. Because of the 90-degree phase shift between the current components, the total current density is only $\sim 0.5\%$ larger than its resistive component, with a small phase difference of around $6\degree$ between them (and between the current density and E-field). Although the modeled capacitive conductance reaches more than half the amplitude of the resistive conductance at frequencies below 100 Hz, these (extrapolated) values are considered conservative and have high uncertainties in the low frequency range due to the limitation of the measurement setup. Further, this low frequency range constitutes a small fraction of the spectrum. Even within the wider range from 0 Hz to 1 MHz, the data by Gabriel et al. indicate the average capacitive-to-resistive conductance ratio is $\sim 20\%$, introducing only a $\sim 2\%$ amplitude error and $\sim 11\degree$ phase difference in the current density. Such small errors may accumulate for high pulse rate monophasic stimulation, in which the capacitive charge does not have time to dissipate between pulses [5], but such accumulation is mitigated by charge balanced biphasic pulses or low pulse repetition rates.
@@
The variations of the resistive conductance of neural tissue are generally considered relatively small over the limited frequency range of a decade for charge-balanced pulsed stimulation with a center frequency above 100 Hz. Per the data by Gabriel et al., for example, the conductivity of gray and white matter increases from 0.09 to 0.13 S/m and from 0.06 to 0.08 S/m, respectively, over three decades (Fig. 2 B), corresponding to a change of $\sim 13\%$ per decade on average [83, 3]. Accordingly, the use of non-dispersive fixed brain tissue conductivity values in most neuromodulation modeling papers to represent the parameters near the center frequency of the pulse is mostly a valid approximation under assumption A4.
Overall, the field errors introduced by QSA are generally considered small and would have similar relative effects on neural activation [3, 5, 6]. For quasi-uniform E-fields in non-invasive stimulation, an $x\%$ relative error in the regional E-field amplitude versus ground truth changes the predicted subthreshold polarization by the same relative amount, and approximately results in an opposite change by $-x\%$ (${(1+x\%)}^{-1}\approx 1-x\%$ for small $x\%$) for the predicted minimum stimulation amplitude (current or voltage) needed to achieve suprathreshold activation in that same region. For non-uniform E-field (e.g., near implanted electrodes), the errors in activation thresholds are typically amplified compared to those of the E-field [88] as polarization of axons is proportional to the derivative of the axial E-field (i.e., second derivative of extracellular potentials along the axon). Nevertheless, these QSA errors are typically smaller than the errors of field amplitude due to uncertainty of tissue conductivity, which can have a wide range of values in the literature with max-to-min ratio between 1.5 to over 10 or $50\%$ and higher deviation compared to the mean [82, 89, 90, 91, 92, 6]. Therefore, the selection of the fixed frequency and of the corresponding conductivity can have profound effects on the modeled E-field and predicted activation thresholds due to the sensitivity of the solutions to the conductivity [3, 6, 5]. Further discussions on the accuracy and validity of QSA in the context of scientific questions are given in sections 3.1 and 4, with application-specific examples given in sections 3.3–3.6.

### v5

@@ Abstract
We define and explain the quasistatic approximation (QSA) as applied to field modeling for electrical and magnetic stimulation. Neuromodulation analysis pipelines include discrete stages, and QSA is applied specifically when calculating the electric and magnetic fields generated in tissues by a given stimulation dose. QSA simplifies the modeling equations to support tractable analysis, enhanced understanding, and computational efficiency. The application of QSA in neuro-modulation is based on four underlying assumptions: (A1) no wave propagation or self-induction in tissue, (A2) linear tissue properties, (A3) purely resistive tissue, and (A4) non-dispersive tissue. As a consequence of these assumptions, each tissue is assigned a fixed conductivity, and the simplified equations (e.g., Laplace’s equation) are solved for the spatial distribution of the field, which is separated from the field’s temporal waveform. Recognizing that electrical tissue properties may be more complex, we explain how QSA can be embedded in parallel or iterative pipelines to model frequency dependence or nonlinearity of conductivity. We survey the history and validity of QSA across specific applications, such as microstimulation, deep brain stimulation, spinal cord stimulation, transcranial electrical stimulation, and transcranial magnetic stimulation. The precise definition and explanation of QSA in neuromodulation are essential for rigor when using QSA models or testing their limits.
@@ 1 Introduction
Computational modeling of electromagnetic fields has broad application in neuromodulation to estimate delivered intensity, predict and optimize response, and inform risk. The quasistatic approximation (QSA) [1, 2, 3, 4, 5, 6, 7], applied under the associated quasistatic assumptions, is ubiquitous in computational modeling of electrical and magnetic stimulation. The purpose of QSA is to reduce the computational complexity of calculating the electric fields (E-fields), potentials, and currents in the target tissue, e.g., brain, spinal cord, or peripheral nerve. Despite its prevalence in neuromodulation modeling, QSA as used in neuromodulation has not been clearly defined, and the underlying assumptions are typically implicit or, at times, ambiguous. Thus, the use of QSA, as understood and applied in neuromodulation modeling, needs to be defined. Here we explain how QSA is applied in a consistent manner across modeling of neuromodulation under four underlying assumptions, and we derive the corresponding governing equations. We also review adaptations of the QSA simulation pipeline to accommodate cases where some of the simplifying assumptions do not hold. We further provide historical and application-specific examples and clarify QSA-related terminology as applied in general electromagnetism. Finally, we provide recommendations for communicating the use of QSA in modeling studies.
@@ 1.1 Scope
The scope of this paper includes the computational methods used to predict the delivery of electromagnetic energy to the body during neuromodulation—and specifically the application of QSA. These computations yield the E-fields in the tissue of interest, and/or related quantities including tissue potentials, current densities, or magnetic fields. The solutions provide estimates of these variables, which are subject to model assumptions and verification. We define these assumptions and related parameters, focusing on the E-field, as the magnetic field is not influenced by the tissues when QSA is valid, and as the E-field is understood to mediate the effects of the most common forms of neural stimulation [8]. Nonetheless, the magnetic field has a key role in neural stimulation through electromagnetic induction, and QSA typically applies to the magnetic field as well; therefore, we occasionally note generalization to the magnetic field. To frame the use of QSA, we also broadly consider additional steps in modeling pipelines, before or following the prediction of the E-field. Quasistatic analysis of approaches for recording bioelectric phenomena, such as the electrocardiogram and impedance plethysmography, was summarized by Plonsey and Heppner [1]. Specific to the nervous system, QSA has been applied to analyses of endogenous brain activity as recorded by electroencephalography (EEG) [9, 10] and magnetoencephalography (MEG) [9, 11]. Subsequently, QSA in neuromodulation was adapted from these analyses of bioelectric signal recording. In contrast to endogenous electromagnetic fields, however, neuromodulation uses exogenous fields and applies a wide range of stimulation waveforms with much broader frequency content and much higher intensities than intrinsic signals, and thus warrants distinct considerations. The application of QSA in modeling neuromodulation encompasses several modalities, including:
@@ 2.1 Multi-stage framework for modeling of neuromodulation
In some modeling pipelines, components of the stimulation device (electrodes or coils) are combined with the tissue when the E-field is computed. This combination preserves QSA as long as all underlying assumptions applied to the tissue also apply to the device components. Alternatively, the influence of the electrodes and their interface with tissue can be modeled in a step after defining the current or voltage source and preceding tissue E-field calculation [21, 22], in which case the electrodes are not subject to QSA. For example, a lumped-circuit model can represent the impedance of the electrode-tissue interface. Similarly, the magnetic field generated by a TMS coil can be computed prior to simulation of the induced E-field in tissue, and can account, for example, for nonlinear effects due to saturation of a ferromagnetic core in the coil [23]. These approaches preserve QSA for the E-field in the tissue, allowing for modeling adaptations (section 2.6) to deal with sources with frequency-dependent or nonlinear effects, such as charge transfer conductances of the electrode–electrolyte interface in electrical stimulation and skin effect in coil windings or saturation of ferromagnetic core materials in magnetic stimulation.
@@ 2.2.3 Purely resistive tissue (A3)
Under assumption A3, the time derivative relating the current density and E-field is neglected. Only the free current density and thus the real part of the complex conductivity is considered. This removes the frequency- and location-dependent temporal phase shift between the E-field and current density distributions due to the tissue capacitance. This assumption is typically valid for dc or low permittivity media at low frequencies, and is commonly denoted as “no capacitance” or a “purely resistive” medium in the neuromodulation modeling literature.
@@ 2.2.4 Non-dispersive tissue (A4)
Assumption A4 considers tissue properties to be non-dispersive, indicating that they remain frequency-independent, at least within the spectrum of the stimulation waveform. In practice, neuromodulation modeling studies typically make this assumption implicitly, along with assumptions A2 and A3, by assigning a single conductivity value to each tissue. For stimulation using dc or a single frequency (e.g., tDCS or tACS, respectively), these tissue parameters reflect the magnitude of the impedance at the given frequency. In these cases, assumption A4 is not so much an assumption of frequency-independent tissue properties as the (implicit) recognition that the relevant properties correspond to a specific frequency. For the majority of cases (e.g., pulsed stimulation), waveforms are composed of multiple frequencies (see section 2.4 and Fig. 2 A), and the use of a single parameter for each tissue reflects an approximation. Such an approximation may be justified by considering the central frequency content of the waveform, or otherwise sufficient given the goals of the stimulation study. In all cases, the selection of single fixed purely-resistive tissue values (which involves assumptions A1, A2, and A3) should be recognized as a key step in model development (see section 4).
@@ 2.3.1 Current density in tissue volume conductor
We start by deriving the equation that governs the electric current density $\boldsymbol{J}$ in the tissue under the generalized physical case, with no approximations applied. The current density, absent magnetizing currents in a non-magnetic volume conductor, consists of the free current and displacement current (including the polarization current) driven by the time-varying E-field, $\boldsymbol{E}(t)$,
@@
Current continuity on the boundary of two tissue domains 1 and 2 [61, 62] results in $\sigma_{1}\cdot\boldsymbol{E_{1,\perp}}=\sigma_{2}\cdot\boldsymbol{E_{2,\perp}}$ for the normal component, or specifically $\boldsymbol{E_{1,\perp}}=0$ if domain 2 is non-conductive, such as air or electrode insulation.
@@ 2.3.2 Electrical stimulation
For electrical stimulation, defining the E-field as the gradient of a scalar potential $\varphi$, $\boldsymbol{E}=-\nabla\varphi$, results in the governing equation for electric potential with the real-valued, fixed, and frequency-independent conductivity, which is commonly used in neuromodulation:
@@ 2.3.3 Magnetic stimulation
where $\boldsymbol{A}$ is the magnetic vector potential that is related to the coil magnetic field $\boldsymbol{B}$ by $\boldsymbol{B}=\nabla\times\boldsymbol{A}$. Here, despite the temporal variation being considered very slow (relative to typical radio frequencies [65]) under QSA, the primary field $\boldsymbol{E_{\mathrm{p}}}=-\frac{\partial\boldsymbol{A}}{\partial t}$ generated by the forward mutual induction between the coil and tissue is included because the rate-of-change is high due to the very high coil currents (on the order of kiloamperes). In contrast, the induced currents in the tissue (less than 1 A) are several orders of magnitude smaller than the kA-level coil currents. Therefore, their magnetic field and induced E-field—both self-induction in the tissue and backward mutual induction to the coil—are negligible. The secondary field $\boldsymbol{E_{\mathrm{s}}}=-\nabla\varphi$ comes from the charge densities on tissue boundaries due to imbalance of the current density of the primary E-field, in a process considered instantaneous under QSA [60]. Given (3), the scalar potential is related to the vector potential by [66, 67, 4, 68]
$\nabla\cdot(\sigma\nabla\varphi)=-\nabla\cdot(\sigma\frac{\partial\boldsymbol{A}}{\partial t})=-\nabla\cdot(\sigma\boldsymbol{E_{\mathrm{p}}})$ (7)
and can be obtained numerically via techniques such as the finite- or boundary-element method (FEM or BEM, respectively).
In magnetic stimulation, the presence of the volume conductor comprising tissues does not significantly affect the electrical characteristics of the coils or their magnetic field. Therefore, the coil field can be pre-calculated or measured in the form of the magnetic field ($\boldsymbol{B}$ or $\boldsymbol{A}$) or the primary field $\boldsymbol{E}_{\mathrm{p}}$ and applied as a background field in the numeric solver [67, 23]. The field solution is computed for either a specific rate-of-change of the coil current (e.g., 1 A/µs) or a value to match the coil current waveform characteristics. Hence, the electromagnetic induction is handled in a stage preceding the quasistatic modeling of the E-field. Alternatively, the coil(s) can be explicitly included as part of the tissue E-field model, and a frequency-domain simulation may be used to calculate the coil magnetic field with the appropriate current amplitude and frequency to reflect the coil current’s rate of change [69].
@@ 2.4 Frequency spectrum of stimulation pulses and frequency dependence of tissue electric parameters
tDCS applies dc current via scalp electrodes (disregarding slow ramp-ups and ramp-downs). Usually, tACS is applied using only a single frequency. When the stimulation waveform is a relatively high frequency sinusoid that is amplitude modulated, the frequency spectrum is centered around the carrier frequency—not around the lower frequency of the amplitude modulation. This is also the case for TIS, where the frequency spectrum consists of the relatively higher frequency of the two (or more) individual channels and does not contain low frequency energy [70, 71, 72].
@@
The frequency spectrum of pulsed stimulation depends on the pulse width, shape, and repetition patterns. While monophasic pulses cover a wide frequency range over a few decades starting from dc (Fig. 2 Ai), neural stimulation almost always uses charge balanced stimulation pulses (Fig. 2 Aii–Aiii) that have energy typically distributed within a frequency range of around one decade. The center location of the spectrum is determined by the pulse duration, typically falling between 100 Hz and 100 kHz. Increasing the total pulse duration–whether by increasing the phase duration, including an interphase interval, or using an asymmetric pulse with a longer and lower amplitude secondary phase–reduces the center frequency of the spectrum but has smaller effects on its shape. The pattern of pulse stimulation (e.g., continuous versus burst) also does not typically change the overall range or shape of the spectrum as governed by the individual pulses, but rather results in a discretized spectrum with the sampling and envelope determined by the repetition pattern and total duration of the pulse train. The spectra of the stimulation waveform should be considered when selecting the conductivity of tissue(s) within neuromodulation models.
@@
Figure 2: A. Normalized power spectrum density between 0 and 100 kHz of representative neurostimulation pulse waveforms, calculated using a 1 µs sampling rate and a total time interval of 1 s for single pulses. i. Monophasic pulses with pulse width (PW) of 1 ms (blue: single pulse; light blue: 30 s pulse train at 20 Hz) and 100 µs (red: single pulse). ii. Charge-balanced biphasic pulses with a first phase of 100 µs duration and with a second phase of equal duration (blue: single pulse; light blue: 15 s pulse train at 125 Hz), a second phase of equal duration and with 100 µs interphase/intrapulse interval (IPI) (red: single pulse), and second phase with 500 µs duration and 20% amplitude of first phase (green: single pulse). iii. E-field waveforms induced by magnetic stimulation pulses, including monophasic (blue: single pulse) and biphasic (red: single pulse; orange: continuous theta burst stimulation (cTBS), 3 pulses per burst, 50 Hz intraburst frequency, 5 Hz interburst frequency, and 40 s duration), recorded from a MagPro X100 device with a Mag-Venture MCF-B70 figure-of-8 coil (MagVenture A/S, Farum, Denmark) [33, 38]. Electromagnetically-induced pulses are automatically charged-balanced as the amplitude of the coil current returns to zero at the end of a pulse. B. Spectra of tissue electrical parameters between 0 and 1 MHz for gray matter (blue) and white matter (red) as modeled by Gabriel et al. [57]. i. Relative permittivity. ii. Resistive (darker colors) and capacitive (lighter colors) conductivities. iii. Ratio between conductive and capacitive conductivities. Conductivity variations are small over the frequency ranges of charge-balanced pulsed stimulation, enabling the approximation using single conductivity values. Dashed lines indicate extrapolation of the model to very low frequencies ($<10$ Hz), outside the range of experimental data (10 Hz to 20 GHz) to which the model was fitted, where the model has unknown validity. Moreover, the measurements had high uncertainties in the low frequency range (10 to 100 Hz) due to the experimental setup. All horizontal axes use a mixed scale, with linear scale between 0 and 10 Hz and logarithmic scale above 10 Hz.
@@
In 1996, Gabriel et al. published a series of papers on the electrical conductivity and permittivity of tissue at frequencies from 10 Hz up to 20 GHz [73, 74, 57]; this series was subsequently updated [75]. Their data and models, which fit particularly well at high frequencies, are broadly consistent with the dielectric properties measured in other studies and as described by the Cole–Cole equation, and have, to some extent, become a canonical reference for modeling studies explicitly considering the suitability of QSA.
@@
According to the data by Gabriel et al., the capacitive current of neural tissues in the relevant frequency range is $\sim 10\%$ compared to the resistive current (Fig.2 B). The total current density consists of the resistive and capacitive currents in parallel and is driven by the same E-field. Because of the 90-degree phase shift between the current components, the total current density is only $\sim 0.5\%$ larger than its resistive component, with a small phase difference of around $6\degree$ between them (and between the current density and E-field). Although the modeled capacitive conductance reaches more than half the amplitude of the resistive conductance at frequencies below 100 Hz, these (extrapolated) values are considered conservative and have high uncertainties in the low frequency range due to the limitation of the measurement setup. Further, this low frequency range constitutes a small fraction of the spectrum. Even within the wider range from dc to 1 MHz, the data by Gabriel et al. indicate the average capacitive-to-resistive conductance ratio is $\sim 20\%$, introducing only a $\sim 2\%$ amplitude error and $\sim 11\degree$ phase difference in the current density. Such small errors may accumulate for high pulse rate monophasic stimulation, in which the capacitive charge does not have time to dissipate between pulses [5], but such accumulation is mitigated by charge balanced biphasic pulses or low pulse repetition rates.
@@
The variations of the resistive conductance of neural tissue are generally considered relatively small over the limited frequency range of a decade for charge-balanced pulsed stimulation with a center frequency above 100 Hz. Per the data by Gabriel et al., for example, the conductivity of gray and white matter increases from 0.09 to 0.13 S/m and from 0.06 to 0.08 S/m, respectively, over three decades (Fig. 2 B), corresponding to a change of $\sim 13\%$ per decade on average [73, 3]. Accordingly, the use of non-dispersive fixed brain tissue conductivity values in most neuromodulation modeling papers to represent the parameters near the center frequency of the pulse is generally considered a valid approximation under assumption A4. However, the selection of the fixed frequency and of the corresponding conductivity can have profound effects on the modeled E-field and predicted activation thresholds, due to the sensitivity of the solutions to the conductivity [3, 5].